OpenStudy (anonymous):

in an orchard there are 150 oranges trees and 250 apple trees . 1.in simplest fraction form what is the ratio of apple trees to orange trees 2.in simplest fraction form what is the ratio of apple trees to orange tress 3.in simplest fraction form what is the ratio of apples trees total trees in the orchard.

OpenStudy (inkyvoyd):

your first two questions are the same

OpenStudy (anonymous):

but it said same on the paper

OpenStudy (inkyvoyd):

are you sure it's tnot apple to orange and orange to apple?

OpenStudy (anonymous):

i am sure

OpenStudy (inkyvoyd):

ok then I did the first two already. for the last one, just take 150/(150+250)=150/400=15/40=3/8 or 3:8 or 3 to 8

OpenStudy (anonymous):

reapted the two first one

OpenStudy (inkyvoyd):

first onee I did already.. it's 3/5 or 3 to 5 or 3:5

OpenStudy (inkyvoyd):

The first two are equal to 3/5. The third one is 3/8

OpenStudy (anonymous):

Ratio of apple tree to orange tree = Apple Tree/Orange tree Ratio = 250/150 Both numbers are divisible by 50 so Ratio = (250/50)/(150/50) = 5/3 Second question have same answer Ratio of apple tree to total trees = Apple Tree/Total trees Ratio = 250/(250 +150) Both numbers are divisible by 50 so Ratio = (250/50)/(400/50) = 5/8